OpenStudy (anonymous):

is 8/13 rational or irrational feel dumb asking this

OpenStudy (anonymous):

rational; any number that can be expressed as a fraction is considered as a fraction

OpenStudy (anonymous):

is it repeating though

OpenStudy (anonymous):

it's repeating

OpenStudy (anonymous):

so irrational?

OpenStudy (anonymous):

no rational

OpenStudy (paxpolaris):

irrationals are NON-repeating NON-terminating decimals

OpenStudy (anonymous):

irrational is a number that can't be written as a ratio such as \[\sqrt{2}\]

OpenStudy (anonymous):

but what about pi its 22/7?

OpenStudy (anonymous):

22/7 is only aproximation of pi

OpenStudy (anonymous):

pi itself is an irrational number

OpenStudy (paxpolaris):

no pi is approximately 22/7

OpenStudy (anonymous):

ok so its rational when it is in that approximated form?

OpenStudy (anonymous):

22/7 is a rational number but it's not actually pi

OpenStudy (anonymous):

a rational number is any number in the form:\[\frac{a}{b}\]where an and b are integers. pi is not a rational number. \[\frac{22}{7}\] is a rational number, because 22 is an integer, and 7 is an integer.

OpenStudy (anonymous):

it doesnt matter if the decimal representation repeats or not.

OpenStudy (anonymous):

ah, thanks guys

OpenStudy (paxpolaris):

\[{22 \over 7} = \begin{matrix} \\ 3.\end{matrix}{\over142857}\] is close enough to pi=3.14159... for everyday use

OpenStudy (zarkon):

\[3.\overline{142857}\] 3.\overline{142857}

OpenStudy (paxpolaris):

any rational number a/b (where b not 0)when written in decimal form will turn out to be terminating or repeating. Any terminating or repeating decimal can be re-written as a/b